OpenStudy (clara1223):

find the limit as x approaches infinity of (4x-7)/((5x^2)+3x+7)

OpenStudy (clara1223):

\[\lim_{x \rightarrow \infty}\frac{ 4x-7 }{ 5x ^{2}+3x+7 }\]

OpenStudy (clara1223):

We haven't gone over limits as x approaches infinity in class before.

OpenStudy (amistre64):

divide it all by the highest power of x to simplify. anything left with an x in the denominator will zero out since 1/inf is a very very small amount of cake to get

OpenStudy (clara1223):

so multiply the denominator by 1/x^2?

OpenStudy (amistre64):

your basically searching for any horizontal asymptotes ... you covered them yet?

OpenStudy (amistre64):

top and bottom yes

OpenStudy (amistre64):

there are 'rules' which you can commit to memory if you have the gigabytes to play with :)

OpenStudy (clara1223):

so I'm left with \[\frac{ \frac{ 4 }{ x }-\frac{ 7 }{ x ^{2} } }{ 5+\frac{ 3 }{ x }+\frac{ 7 }{ x ^{2} } }\]

OpenStudy (amistre64):

now everything with an x under it, goes to something very very small, they approach zero. what are we left with?

OpenStudy (clara1223):

1/5?

OpenStudy (amistre64):

powers of x that is ... 4/x - 7/x^2 is not 1

OpenStudy (clara1223):

-1?

OpenStudy (amistre64):

0-0 = ??

OpenStudy (clara1223):

oh so the answer is 0?

OpenStudy (amistre64):

yep

OpenStudy (clara1223):

because 0/5=0

OpenStudy (amistre64):

one rule, if the bottom is a higher degree than the top, as x approaches infinity, the limit is zero

OpenStudy (amistre64):

if they are of the same degree, they limit to their leading coeffs .. its just simpler to divide off by the highest power of x and assess